A small class has 9 students, 6 of whom are girls and 3 of whom are boys. The teacher is going to choose two of the students at random. What is the probability that the first student chosen will be a boy and the second will be a girl? Write your answer as a fraction in the simplest form.
Answer by math_tutor2020(3816)


A = P(1st is a boy) = (number of boys)/(number of students total)
A = P(1st is a boy) = 3/9
A = P(1st is a boy) = 1/3

B = P(2nd is a girl, given 1st is a boy) = (number of girls)/(number of students left)
B = P(2nd is a girl, given 1st is a boy) = (6)/(9-1)
B = P(2nd is a girl, given 1st is a boy) = 6/8
B = P(2nd is a girl, given 1st is a boy) = 3/4

P(1st is a boy, 2nd is a girl) = A*B = (1/3)*(3/4) = 1/4

Answer: 1/4